OpenStudy (anonymous):

What is the slope of the line whose equation is –48 = 2x − 8y?

jimthompson5910 (jim_thompson5910):

We need to isolate y –48 = 2x − 8y –48-2x = -8y -8y = -2x - 48 y = (-2x - 48)/(-8) y = (-2x)/(-8) - 48/(-8) y = (1/4)x + 6 The last equation above is in the form y = mx+b where m = 1/4 is the slope

jimthompson5910 (jim_thompson5910):

b = 6 is the y-intercept so that graph crosses the y axis at (0,6)

OpenStudy (anonymous):

?

jimthompson5910 (jim_thompson5910):

where are you stuck?

OpenStudy (anonymous):

are u sure its 1/4

OpenStudy (anonymous):

i got 7/4

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

how did you get 7/4 ?

OpenStudy (anonymous):

tbh idk

jimthompson5910 (jim_thompson5910):

do you see how I isolated y?

OpenStudy (anonymous):

i have 2 mor for u is tht ok

jimthompson5910 (jim_thompson5910):

we're not done with this one though.

jimthompson5910 (jim_thompson5910):

do you see how I isolated y? or no?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

so what would be the slope for something like 3y = 10x + 7

OpenStudy (anonymous):

explain plz

jimthompson5910 (jim_thompson5910):

again the idea is to isolate y so what do we do to accomplish this?

OpenStudy (anonymous):

http://openstudy.com/study#/updates/541e7514e4b0a68c6c059e85 help me understand this one instead

OpenStudy (anonymous):

What is the slope of the line whose equation is 12 = 4x – 6y? this is more like the last one